# Pitfalls of Scale: Investigating the Inverse Task of Redefinition in Large Language Models

### v1

@@ Appendix E Results on units of measure redefinition
An overview of response accuracy is presented in Table 18, where we consider the hardest redefinitions corresponding to $R_{a}3$ and $R_{s}2$ for assignment and swapping respectively, as well as all three question levels, together with FF and MC response formats regarding units of measurement redefinitions. Additionally, Table 19 presents the number of correct responses in the NR case alongside the anchoring rate for FF responses regarding units redefinitions. Anchoring is less prominent for some LLMs in comparison to their anchored responses in the constants redefinition task; for instance, Command r+, light text, text achieve even 0% anchoring in some cases, even in $Q_{3}$ questions over the hardest $R_{a}3$ unit redefinitions. Nevertheless, anchoring still persists in many instances, with large rates concerning models in the Mistral family for the hardest question and redefinition levels. Moreover, Titan models present high anchoring even for $R_{a}2$ unit redefinitions, even in the easier $Q_{1}$ level. Surprisingly, anchoring for Titan models reduces as questions and redefinitions become harder, but this does not indicate an improvement in producing correct responses and therefore an advancement in reasoning capability; instead, the anchoring reduction is attributed to the generation of more completely wrong responses, indicating those models’ inability of solving the unit of measure redefinition task appropriately.
@@
Claude sonnet v2 0.0 0.0 0.0 12.5 6.25 6.25 0.0 0.0 0.0 6.25 0.0 0.0
Claude sonnet v2 S 6.25 18.75 6.25 31.25 18.75 31.25 6.25 0.0 6.25 25.0 6.25 12.5
Table 18: The percentage of anchored responses for all LLMs tested under the ZS prompting setup for the most difficult units of measure redefinitions in assignment ($R_{a}3$) and swapping ($R_{a}2$). The highest rate for each model family is presented in bold.
@@
Claude sonnet v1 100.0 0.0 87.5 0.0 87.5 6.25 100.0 0.0 87.5 0.0 87.5 0.0
Claude sonnet v2 93.75 6.25 68.75 6.25 25.0 18.75 93.75 6.25 68.75 6.25 25.0 6.25
@@
Furthermore, Figures 13 and 14 illustrate the results of Mistral7B and Mistral Large, as well as Llama8B and Llama 405B, respectively, regarding units of measure redefinitions. Once again, Mistral7B tends to provide fewer anchored responses compared to its larger counterpart. The same trend holds for Llama, although for $Q_{1}$ and $Q_{2}$ responses, the increase in anchoring is less pronounced for the larger model.
@@ Appendix G Prompts
This section illustrates the prompts used for the LLMs. The prompts vary based on the task (NR or redefinition), the required response format (FF or MC), and the prompting techniques selected (ZS, FS, or CoT).
@@
The answers generated by the LLMs were parsed through a different model (Claude 3.5 Sonnet v2) to determine whether they matched the correct response or the anchored one. This approach was necessary because LLMs can produce additional output (e.g. using CoT), making it difficult to extract their answers using regular expressions. The prompts used in this procedure for the NR and redefinition tasks are presented below.
@@ 4. If the LLM answer matches neither of the two, output: none
Lastly, to categorize the answers as completely wrong, refused, or blank, we prompted Claude 3.5 Sonnet v2 using the following prompt

### v2

@@ Appendix E Results on units of measure redefinition
An overview of response accuracy is presented in Table 18, where we consider the hardest redefinitions corresponding to the $R_{a}2$ and $R_{a}3$ assignment types, as well as all three question levels, together with FF and MC response formats regarding units of measurement redefinitions. Additionally, Table 19 presents the number of correct responses in the NR case alongside the anchoring rate for FF responses regarding units redefinitions. Anchoring is less prominent for some LLMs in comparison to their anchored responses in the constants redefinition task; for instance, Command r+, light text, text achieve even 0% anchoring in some cases, even in $Q_{3}$ questions over the hardest $R_{a}3$ unit redefinitions. Nevertheless, anchoring still persists in many instances, with large rates concerning models in the Mistral family for the hardest question and redefinition levels. Moreover, Titan models present high anchoring even for $R_{a}2$ unit redefinitions, even in the easier $Q_{1}$ level. Surprisingly, anchoring for Titan models reduces as questions and redefinitions become harder, but this does not indicate an improvement in producing correct responses and therefore an advancement in reasoning capability; instead, the anchoring reduction is attributed to the generation of more completely wrong responses, indicating those models’ inability of solving the unit of measure redefinition task appropriately.
@@
Claude v2 6.25 18.75 6.25 31.25 18.75 31.25 6.25 0.0 6.25 25.0 6.25 12.5
Claude 3.5 Sonnet 0.0 0.0 0.0 12.5 6.25 6.25 0.0 0.0 0.0 6.25 0.0 0.0
Claude 3.7 Sonnet 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Table 18: The percentage of anchored responses for all LLMs tested under the ZS prompting setup for the most difficult units of measure redefinitions ($R_{a}2$ and $R_{a}3$ levels). The highest rate for each model family is presented in bold.
@@
Claude v2 93.75 6.25 68.75 6.25 25.0 18.75 93.75 6.25 68.75 6.25 25.0 6.25
Claude 3.5 Sonnet 100.0 0.0 87.5 0.0 87.5 6.25 100.0 0.0 87.5 0.0 87.5 0.0
Claude 3.7 Sonnet 100.0 0.0 87.5 0.0 93.75 0.0 100.0 0.0 87.5 0.0 93.75 0.0
@@
Furthermore, Figures 14 and 15 illustrate the results of Mistral 7B and Mistral Large, as well as Llama 8B and Llama 405B, respectively, regarding units of measure redefinitions. Once again, Mistral 7B tends to provide fewer anchored responses compared to its larger counterpart. The same trend holds for Llama, although for $Q_{1}$ and $Q_{2}$ responses, the increase in anchoring is less pronounced for the larger model.
@@ Appendix F Implementation details
Claude 3.7 Sonnet anthropic/claude-3.7-sonnet https://openrouter.ai/anthropic/claude-3.7-sonnet
@@ Appendix G Prompts
This section illustrates the prompts used to question the LLMs. The prompts vary based on the task (NR or redefinition), the required response format (FF or MC), and the prompting techniques selected (ZS, FS, or CoT).
@@
The answers generated by the LLMs are parsed through a different model (particularly Claude 3.5 Sonnet) to determine whether they match the correct response or the anchored one. This approach is necessary because LLMs can produce additional outputs (e.g. using CoT), making it difficult to extract their answers using regular expressions. The prompts used in this procedure for the NR and redefinition tasks are presented below.
@@ 4. If the LLM answer matches neither of the two, output: none
Lastly, to categorize the answers as completely wrong, refused, or blank, we prompt Claude 3.5 Sonnet using the following prompt.